# Defending Jailbreak Attack in VLMs via Cross-modality Information Detector

### v2

@@ 4.2 Effectiveness
DSR. We first demonstrate the overall DSR that CIDER can achieve and compare it with the baseline method, Jailguard. Table 1 shows that CIDER achieves a DSR of approximately 80%, while the DSR of Jailguard varies, depending on the target VLMs. Note that CIDER is independent of the VLMs, thus the DSR does not vary with the choice of VLMs. However, Jailguard’s detection capability relies heavily on the model’s safety alignment, so the DSR also varies. VLMs with good alignment achieve high DSR (e.g., GPT4V), while poorly aligned VLMs have relatively low DSR (e.g., InstructBLIP). In other words, Jailguard does not significantly enhance VLM robustness against adversarial jailbreak attacks, whereas CIDER does. Nonetheless, CIDER achieves a higher DSR than most of the Jailguard results, except Jailguard on GPT4V.
@@
ASR. To evaluate the effectiveness of CIDER, we measure the decline in ASR after applying CIDER. Figure 5 compares the original ASR without defense (red bar), ASR after CIDER (blue bar) and ASR after Jailguard (yellow bar). Note that, Jailguard is solely designed to detect jailbreak input. To ensure a fair comparison, we add an output module following Jailguard’s detection. Specifically, if Jailguard detects a jailbreak, it will refuse to respond, similar to CIDER. Otherwise, the original input will be processed by the VLM.
Across all models, defending with CIDER significantly reduces the ASR, yielding better results than the baseline. This indicates that CIDER effectively enhances the robustness of VLMs against optimization-based jailbreak attacks. The most notable improvements are seen in LLaVA-v1.5-7B, where ASR drops from 60% to 0%, and in MiniGPT4, from 57% to 9%. For VLMs with initially low ASRs, such as InstructBLIP and Qwen-VL, ASR is reduced to approximately 2% and 1% respectively. Another notable disadvantage of Jailguard is observed in models like GPT4V, InstructBLIP, and Qwen-VL, which already had strong safety alignment and resistance to adversarial attacks. In these cases, the use of Jailguard resulted in a slight increase in ASR.
@@ 4.3 Efficiency
Timely inference is crucial for safeguarding VLMs in real-world applications. Table 2 shows the time required to process a single input pair and generate up to 300 tokens with different VLMs, comparing no defense, CIDER, and Jailguard.
@@ 4.4 Robustness-utility trade-off
Figure 6: VLM performance with and without CIDER on MM-Vet.
To further demonstrate CIDER’s influence on the original utilities on normal queries, we also evaluate the utility of CIDER protected VLMs on MM-Vet benchmark, including recognition, OCR, knowledge, language generation, spatial awareness, and math. As shown in Figure 6, employing CIDER leads to an approximate 30% overall performance decline on normal tasks. Specifically, CIDER mostly affects the VLM’s recognition, knowledge, and language generation capabilities, while it has minimal impact on OCR, spatial awareness, and math skills. We hypothesize that CIDER’s stringent decision-making process, which outright rejects tasks once an image is identified as adversarial, hampers the model’s overall performance. To further illustrate the robustness-utility trade-off, we conducted an ablation study using denoised images as inputs for the adversarial images, termed CIDER-de. The result is relegated to Appendix B.
@@ 4.5 Generalization
In the previous sections, we evaluated the ASR and DSR against adversarial examples generated by Qi et al. (2024). To assess the generalization of our defense method, which is crucial for its applicability to other attacks, we evaluate CIDER against another optimization-based jailbreak attack. We generated 800 adversarial pairs using ImgJP, as proposed by Niu et al. (2024). Table 3 presents the drop in ASRs on four open-source VLMs. The ASR for all VLMs dropped to below 4%, with Qwen reaching 0%. Additionally, CIDER achieves a DSR of 93.87% against ImgJP. These results demonstrate that CIDER effectively generalizes in defending against optimization-based adversarial attacks, highlighting its practical utility for real-world applications.
Base CIDER
Model ASR(%) ASR(%) $\Delta$ (%)
@@
Table 3: Generalization of CIDER to ImgJP
@@ 5 Related Work
Vision Language Model. A typical Vision Language Model (VLM) consists of an image encoder Dosovitskiy et al. (2020) to extract feature maps, a projector to align image modality information with text modality, and a Large Language Model (LLM) to integrate textual and visual input for generating responses. The impressive multimodal capabilities of these models have spurred significant research interest, leading to contributions from both academia and industry Achiam et al. (2023); Liu et al. (2024); Zhu et al. (2023); Dai et al. (2024); Bai et al. (2023).
Jailbreaking VLMs. Incorporating visual information into the LLM framework significantly broadens its range of applications but also introduces new security vulnerabilities, complicating the security issues of VLMs. Besides transferring text jailbreak templates from LLMs to VLMs Luo et al. (2024), effective strategies for jailbreaking VLMs include using gradient-based methods to generate adversarial images Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024), and submitting screenshots containing harmful instructions Gong et al. (2023) or related images Liu et al. (2023); Shayegani et al. (2023). This paper focuses on safeguarding VLMs against gradient-based adversarial image attacks, aiming to fortify VLMs against such sophisticated threats and ensure their robustness and reliability in practical applications.
Safeguarding VLMs. Various defense mechanisms have been proposed to address vulnerabilities in VLMs and enhance their security and robustness. These mechanisms can be categorized into proactive and reactive defenses based on their preventive and responsive nature. Proactive defenses aim to prevent attacks through techniques like adversarial training Zong et al. (2024) and reinforcement learning Chen et al. (2023) during the training phase. In contrast, reactive defenses focus on detecting attacks during the inference phase using methods such as Wang et al. (2024a); Pi et al. (2024); Wang et al. (2024b). However, many of these methods require access to internal model parameters or rely on additional large models for implementation. Our approach prioritizes a reactive defense strategy for its practicality and ease of implementation. Notably, Jailguard Zhang et al. (2023) is closely related to our work, as it detects jailbreak queries by analyzing variations in responses to perturbed inputs. However, Jailguard’s detection success heavily depends on the safety of the underlying LLM and involves significant computational costs.
@@ 6 Conclusion
In this work, we propose a plug-and-play cross-modality information detector, CIDER, which can effectively and efficiently defend against adversarial jailbreak attacks. Compared to previous methods, CIDER achieves superior defense performance, as evidenced by higher DSR and a significant decline in ASR, while greatly reducing processing time. We also evaluate the transferability of CIDER to other optimization-based adversarial attacks and demonstrate the robustness-utility trade-off in VLMs. In future research, we aim to improve CIDER by reducing the negative impact on VLM utilities to normal tasks. Additionally, it would be useful to develop defense mechanisms against non-optimization-based jailbreak attacks.
@@
1. While CIDER is an effective, efficient, and user-friendly defense mechanism, it does impact VLM performance to some extent. We believe this is due to CIDER’s stringent handling of adversarial examples. In future work, we plan to implement multi-level thresholds to process adversarial examples with varying degrees of rigor, aiming to maintain robust defense capabilities without compromising VLM performance.
2. CIDER is specifically designed to defend against optimization-based adversarial jailbreak attacks, and its effectiveness against other types of jailbreak attacks is uncertain. Future research will explore CIDER’s effectiveness against these alternative attacks and develop corresponding defense strategies, aiming to enhance the overall security and resilience of VLMs against a wider array of adversarial threats.
@@
Ensuring the security of Vision Large Language Models (VLMs) is crucial as they become more widely used in various applications. This paper introduces CIDER, a simple yet effective cross-modality information detector designed to defend against adversarial jailbreak attacks in VLMs. Our work significantly contributes to the field by providing a tool that mitigates known vulnerabilities and lays the groundwork for future improvements in safety measures. While CIDER marks significant progress, it doesn’t make VLMs immune to all threats. Continuous evaluation and updates are crucial as VLMs evolve. By sharing CIDER and our findings, we aim to encourage ongoing research and collaboration, promoting advanced and secure VLMs. We are committed to addressing the ethical implications of VLM deployment, ensuring transparency, and prioritizing the responsible use of these technologies for societal benefit.
@@ Appendix B Ablation study on robustness-utility trade-off
To further illustrate the robustness-utility trade-off, we perform an ablation study using denoised images as inputs for adversarial images, referred to as CIDER-de. Figure 8 shows the ASR of CIDER-de and Figure 9 shows the MM-Vet score of it. It can be observed that using CIDER-de hardly impacts the utility of the VLM. However, this comes at the expense of greatly diminished defensive effectiveness.
Figure 8: ASR of base VLM, defending with CIDER-de, CIDER and Jailguard
Figure 9: MM-Vet score of base VLM, defending with CIDER-de, CIDER and Jailguard

### v4

@@ 4.2 Effectiveness
DSR. We first demonstrate the overall DSR that CIDER can achieve and compare it with the baseline method, Jailguard. Table 1 shows that CIDER achieves a DSR of approximately 80%, while the DSR of Jailguard varies, depending on the target MLLMs. Note that CIDER is independent of the MLLMs, thus the DSR does not vary with the choice of MLLMs. However, Jailguard’s detection capability relies heavily on the model’s safety alignment, so the DSR also varies. MLLMs with good alignment achieve high DSR (e.g., GPT4V), while poorly aligned MLLMs have relatively low DSR (e.g., InstructBLIP). In other words, Jailguard does not significantly enhance MLLM robustness against adversarial jailbreak attacks, whereas CIDER does. Nonetheless, CIDER achieves a higher DSR than most of the Jailguard results, except Jailguard on GPT4V.
@@
ASR. To evaluate the effectiveness of CIDER, we measure the decline in ASR after applying CIDER. Figure 5 compares the original ASR without defense (red bar), ASR after CIDER (blue bar), and ASR after Jailguard (yellow bar). Note that, Jailguard is solely designed to detect jailbreak input. To ensure a fair comparison, we add an output module following Jailguard’s detection. Specifically, if Jailguard detects a jailbreak, it will refuse to respond, similar to CIDER. Otherwise, the MLLM will process the original input.
Across all models, defending with CIDER significantly reduces the ASR, yielding better results than the baseline. This indicates that CIDER effectively enhances the robustness of MLLMs against optimization-based jailbreak attacks. The most notable improvements are seen in LLaVA-v1.5-7B, where ASR drops from 60% to 0%, and in MiniGPT4, from 57% to 9%. For MLLMs with initially low ASRs, such as InstructBLIP and Qwen-VL, ASR is reduced to approximately 2% and 1% respectively. Another notable disadvantage of Jailguard is observed in models like GPT4V, InstructBLIP, and Qwen-VL, which already had strong safety alignment and resistance to adversarial attacks. In these cases, the use of Jailguard resulted in a slight increase in ASR.
We conclude that the threshold determined by CIDER can be effectively applied to different MLLMs due to their shared transformer-based LLM backbones, which generate comparable representations of harmful information. This harmful information, distilled from malicious queries, is embedded into adversarial images using similar optimization-based attacks. As a result, the consistent noise patterns produced by these attacks across different MLLMs can be detected using the same threshold, highlighting the robustness and transferability of CIDER.
@@ 4.3 Efficiency
Timely inference is crucial for safeguarding MLLMs in real-world applications. Table 2 shows the time required to process a single input pair and generate up to 300 tokens with different MLLMs, comparing no defense, CIDER, and Jailguard.
@@ 4.4 Robustness-utility trade-off
Figure 6: MLLM performance with and without CIDER on MM-Vet.
To further demonstrate CIDER’s influence on the original utilities on normal queries, we also evaluate the utility of CIDER protected MLLMs on MM-Vet benchmark, including recognition, OCR, knowledge, language generation, spatial awareness, and math. As shown in Figure 6, employing CIDER leads to an approximate 30% overall performance decline on normal tasks. Specifically, CIDER mostly affects the MLLM’s recognition, knowledge, and language generation capabilities, while it has minimal impact on OCR, spatial awareness, and math skills. We hypothesize that CIDER’s stringent decision-making process, which outright rejects tasks once an image is identified as adversarial, hampers the model’s overall performance. To further illustrate the robustness-utility trade-off, we conducted an ablation study using denoised images as inputs for the adversarial images, termed CIDER-de. The result is relegated to Appendix B.
@@
4.5 Ablation study on denoising method
We perform an ablation study on the choice of denoising method in the CIDER architecture, as it significantly impacts both defense effectiveness and general task performance. Image smoothing methods commonly fall into two categories: DNN-based approaches, such as diffusion-based denoisers Nichol and Dhariwal (2021), and traditional filtering methods, like the Non-local Means filter (NLM; Buades et al., 2011). We compare the ASR and MM-Vet scores of these two denoising methods, with the results presented in Table 3. The NLM filter performs similarly to the diffusion-based denoiser in terms of ASR, except on the LLaVA model, where it falls significantly behind. However, for general tasks, the NLM filter underperforms the diffusion-based denoiser across all models. This suggests that while the diffusion-based denoiser effectively reduces noise, it also preserves essential image details, making it a superior choice overall.
[…]
Denoiser NLM Denoiser NLM
LLaVA-v1.5-7B $0.00$ $8.12$ $19.1$ $18.9$
InstructBLIP $2.34$ $2.56$ $12.7$ $11.1$
MiniGPT4 $9.22$ $9.06$ $13.0$ $11.2$
Qwen-VL $1.09$ $0.78$ $29.2$ $27.6$
Average 3.16 $5.13$ 18.5 $17.2$
Table 3: The ASR and MM-Vet score of different denoising methods.
@@ 4.6 Generalization
In the previous sections, we evaluated the DSR and ASR against adversarial examples generated by Qi et al. (2024). To further assess the generalization of our defense method, which is critical for its applicability to other types of attacks, we test CIDER under different attack settings.
Dataset. In addition to Harmbench, we employ RedTeam-2k Luo et al. (2024) as a harmful query dataset, which compiles harmful queries from five different datasets and filters them according to the safety policies of OpenAI and Llama2, resulting in 2,000 red-teaming questions. We randomly select 200 queries and generate 800 text-image pairs using a similar processing method described in Section 2.2. On this dataset, CIDER achieves a DSR of 81.37%, with the reduction in ASRs presented in Table 4.
Model Base CIDER
ASR(%) ASR(%) $\Delta$ (%)
LLaVA-v1.5-7B $29.87$ $1.87$ $28.00$
InstructBLIP $24.13$ $4.37$ $19.76$
MiniGPT4 $43.75$ $15.63$ $28.12$
Qwen-VL $10.62$ $2.37$ $8.25$
Table 4: Generalization to RedTeam-2k dataset
Attack Method. To evaluate the generalization of our defense against different attack methods, we generated 800 adversarial pairs using ImgJP, an optimization-based approach proposed by Niu et al. (2024). Table 5 shows the decrease in ASRs across four open-source MLLMs, with all ASRs falling below 4%, and Qwen-VL achieving a 0% ASR. Moreover, CIDER attained a DSR of 93.87% against ImgJP. These results demonstrate that CIDER effectively generalizes to defend against optimization-based adversarial attacks, underscoring its practical value in real-world applications.
Model Base CIDER
ASR(%) ASR(%) $\Delta$ (%)
@@
Table 5: Generalization to ImgJP attack
@@ 5 Related Work
Multimodal Large Language Model. A typical Multimodal Large Language Model (MLLM) consists of an image encoder Dosovitskiy et al. (2020) to extract feature maps, a projector to align image modality information with text modality, and a Large Language Model (LLM) to integrate textual and visual input for generating responses. The impressive multimodal capabilities of these models have spurred significant research interest, leading to contributions from both academia and industry Achiam et al. (2023); Liu et al. (2024); Zhu et al. (2023); Dai et al. (2024); Bai et al. (2023).
Jailbreaking MLLMs. Incorporating visual information into the LLM framework significantly broadens its range of applications but also introduces new security vulnerabilities, complicating the security issues of MLLMs. Besides transferring text jailbreak templates from LLMs to MLLMs Luo et al. (2024), effective strategies for jailbreaking MLLMs include using gradient-based methods to generate adversarial images Carlini et al. (2024); Zhao et al. (2024); Qi et al. (2024); Niu et al. (2024), and submitting screenshots containing harmful instructions Gong et al. (2023) or related images Liu et al. (2023); Shayegani et al. (2023). This paper focuses on safeguarding MLLMs against gradient-based jailbreak attacks via adversarial images, aiming to fortify MLLMs against such sophisticated threats and ensure their robustness and reliability in practical applications.
Safeguarding MLLMs. Various defense mechanisms have been proposed to address vulnerabilities in MLLMs and enhance their security and robustness. These mechanisms can be categorized into proactive and reactive defenses based on their preventive and responsive nature. Proactive defenses aim to prevent attacks through techniques like adversarial training Zong et al. (2024) and reinforcement learning Chen et al. (2023) during the training phase. In contrast, reactive defenses focus on detecting attacks during the inference phase using methods such as Wang et al. (2024a); Pi et al. (2024); Wang et al. (2024b). However, many of these methods require access to internal model parameters or rely on additional large models for implementation. Our approach prioritizes a reactive defense strategy for its practicality and ease of implementation. Notably, Jailguard Zhang et al. (2023) is closely related to our work, as it detects jailbreak queries by analyzing variations in responses to perturbed inputs. However, Jailguard’s detection success heavily depends on the safety of the underlying LLM and involves significant computational costs.
@@ 6 Conclusion
In this work, we propose a plug-and-play cross-modality information detector, CIDER, which can effectively and efficiently defend against adversarial jailbreak attacks. Compared to previous methods, CIDER achieves superior defense performance, as evidenced by higher DSR and a significant decline in ASR, while greatly reducing processing time. We also evaluate the generalization of CIDER to other datasets and optimization-based adversarial attacks, and demonstrate the robustness-utility trade-off in MLLMs. In future research, we aim to improve CIDER by reducing the negative impact on MLLM utilities to normal tasks. Additionally, it would be useful to develop defense mechanisms against non-optimization-based jailbreak attacks.
Acknowledgement
We thank all reviewers for their constructive comments. This work is supported by the Shanghai Engineering Research Center of Intelligent Vision and Imaging and the Open Research Fund of The State Key Laboratory of Blockchain and Data Security, Zhejiang University.
@@ Acknowledgement
1. While CIDER is an effective, efficient, and user-friendly defense mechanism, it does impact MLLM performance to some extent. We believe this is due to CIDER’s stringent handling of adversarial examples. In future work, we plan to implement multi-level thresholds to process adversarial examples with varying degrees of rigor, aiming to maintain robust defense capabilities without compromising MLLM performance.
2. CIDER is specifically designed to defend against optimization-based adversarial jailbreak attacks, and its effectiveness against other types of jailbreak attacks is uncertain. Future research will explore CIDER’s effectiveness against these alternative attacks and develop corresponding defense strategies, aiming to enhance the overall security and resilience of MLLMs against a wider array of adversarial threats.
@@
Ensuring the security of Vision Large Language Models (MLLMs) is crucial as they become more widely used in various applications. This paper introduces CIDER, a simple yet effective cross-modality information detector designed to defend against adversarial jailbreak attacks in MLLMs. Our work significantly contributes to the field by providing a tool that mitigates known vulnerabilities and lays the groundwork for future improvements in safety measures. While CIDER marks significant progress, it doesn’t make MLLMs immune to all threats. Continuous evaluation and updates are crucial as MLLMs evolve. By sharing CIDER and our findings, we aim to encourage ongoing research and collaboration, promoting advanced and secure MLLMs. We are committed to addressing the ethical implications of MLLM deployment, ensuring transparency, and prioritizing the responsible use of these technologies for societal benefit.
@@ Appendix B Ablation study on robustness-utility trade-off
To further illustrate the robustness-utility trade-off, we perform an ablation study using denoised images as inputs for adversarial images, referred to as CIDER-de. Figure 8 shows the ASR of CIDER-de and Figure 9 shows the MM-Vet score of it. It can be observed that using CIDER-de hardly impacts the utility of the MLLM. However, this comes at the expense of greatly diminished defensive effectiveness.
Figure 8: ASR of base MLLM, defending with CIDER-de, CIDER and Jailguard
Figure 9: MM-Vet score of base MLLM, defending with CIDER-de and CIDER